# CRISP-NAM: Competing Risks Interpretable Survival Prediction with Neural Additive Models

### v3

@@ Abstract
Competing risks are crucial considerations in survival modelling, particularly in healthcare domains where patients may experience multiple distinct event types. We propose CRISP-NAM (Competing Risks Interpretable Survival Prediction with Neural Additive Models), an interpretable neural additive model for competing risks survival analysis which extends the neural additive architecture to model cause-specific hazards while preserving feature-level interpretability. Each feature contributes independently to risk estimation through dedicated neural networks, allowing for visualization of complex non-linear relationships between covariates and each competing risk. CRISP-NAM demonstrates competitive performance on multiple datasets compared to existing approaches. Code is available at https://github.com/VectorInstitute/crisp-nam
@@ 1 Introduction
Survival analysis provides a robust framework for time-to-event prediction across multiple disciplines including medicine, finance, and manufacturing. Traditionally rooted in statistical methods, this analytical approach focuses on using available covariates to predict when specific events will occur. As an example, in healthcare settings, clinicians leverage patient data such as laboratory values, medical history, and comorbidities to forecast hospital readmission timing for specific conditions. Financial institutions similarly could apply these techniques to customer transaction histories and demographic profiles to assess default probability on credit products.
@@ 2.1 Cox Proportional Hazards Model
Historically, the Cox Proportional Hazards (Cox PH model) model has been a popular choice for survival analysis. It is a semi-parametric, linear model that relates covariates to the hazard function, which characterizes the instantaneous risk of an event occurring at time $t$, given survival up to that time. The Cox PH model model assumes that covariates have a multiplicative effect on the hazard and that their effects are constant over time (proportional hazards assumption). Mathematically, the hazard function under this model is expressed as:
@@
Despite its widespread use and interpretability, the Cox PH model has several limitations. For example, nonlinear relationships must be manually specified must be manually specified (e.g., using splines), or alternatively introduced using nonlinear kernels Cai et al. (2011) which can be challenging in high-dimensional settings. Additionally, this model assumes that the effect of each covariate on the hazard is constant over time. Violations of this assumption can lead to biased estimates. Finally, in scenarios with many features or complex interactions, prior feature engineering or dimensionality reduction may be necessary to avoid convergence issues or unstable estimates.
@@ 2.2 Deep Learning for Survival Analysis
Neural Additive Models (NAMs) Agarwal et al. (2021), a neural-network extension of Generalized Additive Models (GAMs) have been used in machine learning based survival models, examples of which are SurvNAM Utkin et al. (2022) and CoxNAM Xu and Guo (2023). While both SurvNAM and CoxNAM employ NAMs Agarwal et al. (2021) to enhance interpretability in survival analysis, they differ fundamentally in purpose and integration. CoxNAM is a fully trainable survival model that embeds NAMs directly within the Cox proportional hazards framework, enabling inherently interpretable, end-to-end learning of nonlinear feature effects from survival data. In contrast, SurvNAM is a post-hoc explanation method that approximates the predictions of a pre-trained black-box survival model such as a Random Survival Forest (RSF) by fitting a GAM-extended Cox PH model using NAMs as surrogate learners.
It is to be noted that none of the survival models discussed thus far are capable of modelling competing risks, which will be covered next.
@@ 2.3 Competing Risks in Survival Analysis
While conventional survival models primarily address single outcomes, real-world scenarios frequently involve competing risks that fundamentally alter the probability distribution of the primary event. For instance, if a patient dies, the possibility of experiencing a subsequent heart‑related complication is removed, exemplifying a typical competing‑events situation. Two principal methodological frameworks have emerged for analyzing competing risks:
@@
where $T$ represents the event time and $E\in\{1,2,\ldots,K\}$ denotes the event type.
@@
$R_{k}^{sub}(t)=\{i:T_{i}\geq t\text{ or }(T_{i}<t\text{ and }E_{i}\neq k)\}$ (3)
This risk set includes subjects who have either not experienced any event by time $t$ or have experienced a competing event (not event $k$) before time $t$.
@@
$\small\lambda_{k}^{sub}(t|\mathbf{x})=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t,E=k|i\in R_{k}^{sub}(t),\mathbf{x})}{\Delta t}$ (4)
@@ 2.4 Deep Survival Models for Competing Risks
Several deep learning approaches have been proposed to address competing risks in survival analysis. DeepHit Lee et al. (2018) is a joint model for survival analysis with competing risks. It uses a shared representation network followed by cause-specific sub-networks to model the joint distribution of the event time and event type. The model is trained using a combination of the negative log-likelihood and a ranking loss to encourage concordance between predicted risks and observed outcomes. Neural Fine Gray Jeanselme et al. (2023) extends the Fine-Gray sub-distribution model using neural networks to capture non-linear relationships between covariates and sub-distribution hazards. It allows for flexible modelling of competing risks while maintaining the ability to directly estimate cumulative incidence functions.
Despite these advances, a key limitation of existing deep survival approaches for competing risks is their lack of interpretability, especially at the feature level, making it difficult to understand how individual features contribute to risk predictions for different competing events.
To this end, an extension of the Neural Additive Model for competing risks settings is proposed in this paper, resulting in an inherently interpretable survival model. Our approach retains the interpretability and feature-wise transparency of NAMs, while allowing for flexible, non-linear modelling of cause-specific or sub-distribution hazards in competing risks scenarios.
@@ 3.2 Risk-Specific Projections
For each feature $i$ and competing risk $k$, a separate linear projection transforms the feature representation to its contribution to the log-hazard ratio:
$g_{i,k}(\mathbf{h}_{i})=\mathbf{w}_{i,k}^{T}\mathbf{h}_{i}$ (7)
where $\mathbf{w}_{i,k}\in\mathbb{R}^{d}$ is the projection vector for feature $i$ and risk $k$. This risk-specific projection enables the model to capture how the same feature might influence different competing risks differently.
@@ 3.5 Partial Likelihood Loss Function
To train the model, the standard Cox partial likelihood approach, adapted for competing risks Prentice et al. (1978), is implemented. For a dataset with $n$ subjects, the negative log partial likelihood for event type $k$ is
$\small\mathcal{L}_{k}=-\sum_{i:E_{i}=k}\Bigg{[}\eta_{k}(\mathbf{x}_{i})-\log\left(\sum_{j:T_{j}\geq T_{i}}\exp(\eta_{k}(\mathbf{x}_{j}))\right)\Bigg{]}$ (10)
where $E_{i}$ is the event type for subject $i$ and $T_{i}$ is their event or censoring time. The risk set for each event consists of subjects who have not experienced any event by time $T_{i}$, which is the key characteristic of the cause-specific hazards approach.
The overall loss is the sum of the negative log partial likelihoods across all event types:
@@
where $\gamma$ is the $L_{2}$ regularization parameter and $\Theta$ represents all model parameters.
Since many real-world problems involving competing risks suffer from class imbalance, we adopt a risk-frequency-weighted version of the partial likelihood. Specifically, we define:
$\small\mathcal{L}_{k,\omega}=-\omega_{k}\sum_{i:E_{i}=k}\bigg{[}\eta_{k}(\mathbf{x}_{i})-\log\Big{(}\sum_{j:T_{j}\geq T_{i}}\exp\big{(}\eta_{k}(\mathbf{x}_{j})\big{)}\Big{)}\bigg{]}$ (12)
@@
$\mathcal{L}_{\text{weighted}}=\sum_{k}\mathcal{L}_{k,\omega}$ (13)
@@ 3.6 Baseline Hazard Estimation
In the proportional hazards formulation of the subdistribution hazard model, the hazard for event type $k$ is expressed as
$\lambda_{k}(t\mid\mathbf{x})=\lambda_{0k}(t)\cdot\exp\big{(}\eta_{k}(\mathbf{x})\big{)}$ (14)
where $\lambda_{0k}(t)$ is the baseline subdistribution hazard function for event type $k$, and $\eta_{k}(\mathbf{x})$ is the covariate-dependent log-risk function produced by the model.
We do not directly parameterize $\lambda_{0k}(t)$; instead, we estimate the corresponding baseline cumulative hazard function $\Lambda_{0k}(t)$ after training using the Breslow estimator Breslow (1975):
$\small\hat{\Lambda}_{0k}(t)=\sum_{i:T_{i}\leq t,E_{i}=k}\frac{1}{\sum_{j:T_{j}\geq T_{i}}\exp\left(\eta_{k}(\mathbf{x}_{j})\right)}$ (15)
This cumulative baseline hazard can then be used to compute the cumulative incidence function (CIF) for each event type $k$ at test time.
Estimating baseline hazards is required in the proposed approach for several reasons. First, while the neural additive component of CRISP-NAM efficiently learns relative risks between subjects (hazard ratios), the baseline hazard estimation enables translation of these relative measures into absolute risk predictions. This is necessary for clinical decision-making, where understanding the actual probability of an event is more informative than just knowing if one patient has higher risk than another. Second, in competing risks settings, accurate baseline hazard estimation is required for proper calculation of cumulative incidence functions, as shown in Eqs. (16) and (17). Third, the baseline hazard captures the underlying temporal pattern of risk independent of covariates, allowing CRISP-NAM to generate time-dependent predictions at clinically relevant horizons (e.g., 1-year, 5-year risks). Finally, proper evaluation metrics like Brier scores and TD-AUC values at specific time points depend on accurate absolute risk estimation.
@@ 3.7 Prediction of Absolute Risks
To predict the cumulative incidence function (CIF) for each competing event, we combine the cause-specific hazards approach with the relationship between cause-specific hazards and the CIF. For a subject with covariates $\mathbf{x}$, the CIF for event type $k$ at time $t$ is calculated as
[…]
where $S(t|\mathbf{x})$ is the overall survival function representing the probability of not experiencing any event by time $t$, $u$ is the integration variable representing all potential event times between 0 and $t$, $\lambda_{k}$ is the baseline hazard and $S(u-|\mathbf{x})$ denotes the left limit of the survival function (i.e., the probability of not experiencing any event just before time $u$). Here, $t$ represents the fixed time horizon of interest (e.g., 1-year or 5-year risk), while $u$ captures the continuous nature of when events might occur within that horizon. The overall survival function is defined as
[…]
In practice, a discrete approximation by dividing the time interval into small segments and summing the contributions across these segments using the estimated baseline hazards and the predicted risk scores is used to compute these integrals. Concretely, for discrete time points $\{t_{1},t_{2},\dots,t_{M}\}$, the CIF is approximated by
$\hat{F}_{k}(t|\mathbf{x})\approx\sum_{t_{m}\leq t}\hat{S}(t_{m-1}|\mathbf{x})\cdot\hat{\lambda}_{k}(t_{m}|\mathbf{x})$ (18)
@@
$\hat{S}(t_{m-1}|\mathbf{x})=\exp\left(-\sum_{k^{\prime}=1}^{K}\sum_{t_{\ell}<t_{m}}\hat{\lambda}_{k^{\prime}}(t_{\ell}|\mathbf{x})\right)$
and $\hat{\lambda}_{0k}(t_{m})$ is the estimated baseline hazard at time $t_{m}$ for event type $k$.

### v5

@@ Abstract
Competing risks are crucial considerations in survival modelling, particularly in healthcare domains where patients may experience multiple distinct event types. We propose CRISP-NAM (Competing Risks Interpretable Survival Prediction with Neural Additive Models), an interpretable neural additive model for competing risks survival analysis which extends the neural additive architecture to model cause-specific hazards while preserving feature-level interpretability. Each feature contributes independently to risk estimation through dedicated neural networks, allowing for visualization of complex non-linear relationships between covariates and each competing risk. CRISP-NAM demonstrates competitive performance on multiple datasets compared to existing approaches.
@@ 1 Introduction
Survival analysis provides a robust framework for time-to-event prediction across multiple disciplines including medicine, finance, and manufacturing. Traditionally rooted in statistical methods, this analytical approach focuses on using available covariates to predict when specific events will occur. As an example, in healthcare settings, clinicians leverage patient data such as laboratory values, medical history, and comorbidities to forecast hospital readmission timing for specific medical conditions.
Many real-world survival scenarios often involve competing risks where multiple mutually exclusive events can occur. While recent deep learning advances have improved predictive performance in competing risks settings, these models operate as “black boxes”. This opacity poses significant challenges in high-stakes domains like healthcare, where understanding feature contribution to a model’s predictions is crucial for clinical decision-making, regulatory compliance, and building trust with practitioners.
Moreover, multiple jurisdictions have established interpretability requirements across general and sector-specific regulations. For example, Canada’s Artificial Intelligence and Data Act (AIDA) (Government of Canada, 2024) emphasizes risk-based governance with interpretability assessments during development phases. Sector-specific regulations include Food and Drug Administration (FDA) guidance (U.S. Food and Drug Administration, 2024) requiring clear documentation of AI/ML medical device decision-making and Article 13 of the EU AI Act (art, 2024) which requires high risk AI systems to be designed with sufficient transparency and technical capabilities to explain their outputs, accompanied by clear instructions that enable users to properly interpret and appropriately use the system.
In this work, we bridge a critical gap between interpretability and competing risks modelling in deep survival analysis for healthcare applications by introducing CRISP-NAM (Competing Risks Interpretable Survival Prediction with Neural Additive Models). The main contributions of our work are as follows:
  • We extend Neural Additive Models (NAMs) to competing risks settings, whereas existing NAM-based survival models only handle single-event outcomes.
  • We introduce separate projection functions for each risk-feature pair, allowing features to have differential effects across competing events while maintaining interpretability.
  • The proposed model jointly learns all cause-specific hazards in a single model, rather than treating competing events as censoring or requiring separate models.
  • Our model can be used to generate shape functions and feature importance rankings for each competing risk and this would allow practitioners to understand how different covariates influence each outcome.
  • We incorporate risk-frequency weightings to address class imbalances in competing events which is a common challenge that appears in real-world medical datasets.
The next section reviews the relevant background and related work that motivates our proposed approach.
@@ 2.1 Cox Proportional Hazards Model
Historically, the Cox Proportional Hazards (Cox PH) model (Cox, 1972) has been a popular choice for survival analysis. It is a semi-parametric, linear model that relates covariates to the hazard function, which characterizes the instantaneous risk of an event occurring at time $t$, given survival up to that time. The Cox PH model assumes that covariates have a multiplicative effect on the hazard and that their effects are constant over time (proportional hazards assumption). Mathematically, the hazard function under this model is expressed as:
@@
Despite its widespread use and interpretability, the Cox PH model has several limitations. For example, nonlinear relationships must be manually specified (e.g., using splines), or alternatively introduced using nonlinear kernels Cai et al. (2011) which can be challenging in high-dimensional settings. Additionally, this model assumes that the effect of each covariate on the hazard is constant over time. Violations of this assumption can lead to biased estimates. Finally, in scenarios with many features or complex interactions, prior feature engineering or dimensionality reduction may be necessary to avoid convergence issues or unstable estimates.
@@ 2.2 Deep Learning for Survival Analysis
Neural Additive Models (NAMs) Agarwal et al. (2021), a neural-network extension of Generalized Additive Models (GAMs) have been used in machine learning based survival models, examples of which are SurvNAM Utkin et al. (2022) and CoxNAM Xu and Guo (2023). While both SurvNAM and CoxNAM employ NAMs Agarwal et al. (2021) to enhance interpretability in survival analysis, they differ fundamentally in purpose and integration. CoxNAM is a fully trainable survival model that embeds NAMs directly within the Cox proportional hazards framework, enabling inherently interpretable, end-to-end learning of nonlinear feature effects from survival data. In contrast, SurvNAM is a post-hoc explanation method that approximates the predictions of a pre-trained black-box survival model such as a Random Survival Forest (Ishwaran et al., 2008) by fitting a GAM-extended Cox PH model using NAMs as surrogate learners.
Notably, none of the survival models discussed thus far are capable of modelling competing risks, which will be covered next.
@@ 2.3 Competing Risks in Survival Analysis
While conventional survival models primarily address single outcomes, many real-world scenarios involve competing risks that fundamentally alter the probability distribution of the primary event. For instance, if a patient dies, the possibility of experiencing a subsequent heart‑related complication is removed, illustrating a typical competing‑events situation. Two methodological frameworks have emerged for analyzing competing risks:
@@
where $\mathbf{x}$ represents the covariates, $T$ represents the event time and $E\in\{1,2,\ldots,K\}$ denotes the event type.
@@
$R_{k}^{sub}(t)=\{j:T_{j}\geq t\text{ or }(T_{j}<t\text{ and }E_{j}\neq k)\}$ (3)
This risk set includes subjects, $j$, who have either not experienced any event by time $t$ or have experienced a competing event (not event $k$) before time $t$.
@@
$\small\lambda_{k}^{sub}(t|\mathbf{x})=\lim_{\Delta t\to 0}\frac{P(t\leq T<t+\Delta t,E=k|j\in R_{k}^{sub}(t),\mathbf{x})}{\Delta t}$ (4)
@@ 2.4 Deep Survival Models for Competing Risks
Deep Survival Models leverage deep learning techniques to address competing risks in survival analysis, offering the ability to model complex non-linear patterns in risk prediction. DeepHit Lee et al. (2018) is a joint model for survival analysis with competing risks. It uses a shared representation network followed by cause-specific sub-networks to model the joint distribution of the event time and event type. The model is trained using a combination of the negative log-likelihood and a ranking loss to encourage concordance between predicted risks and observed outcomes. Neural Fine Gray Jeanselme et al. (2023) extends the Fine-Gray sub-distribution model using neural networks to capture non-linear relationships between covariates and sub-distribution hazards. It allows for flexible modelling of competing risks while maintaining the ability to directly estimate cumulative incidence functions. Despite these advances, a key limitation of existing deep survival approaches for competing risks is their lack of interpretability, especially at the feature level, making it difficult to understand how individual features contribute to risk predictions for different competing events. In Fig. 1, we depict the current gap in the literature of deep survival models. Specifically, we are interested in these criteria: Interpretability, Non-Linear Modelling and Competing Risks Capability. Models such as SurvNAM and CoxNAM are interpretable, however, they have not been reported to be used in competing risks settings. In contrast, the Neural Fine-Gray and DeepHit architectures can model competing risks, but are “black-box” models and can only be explained using post-hoc explainability methods such as SHAP and Partial Dependence Plots. Post hoc methods are known to be imprecise and can generate misleading explanations Huang and Marques-Silva (2024); Laugel et al. (2019). The original formulations of Fine-Gray and Cox PH models are incapable of modelling non-linear relationships. Our proposed CRISP-NAM model addresses these gaps in survival models fulfilling all 3 criteria while providing competitive performance.
Figure 1: Venn diagram situating CRISP-NAM against existing deep survival methods
To this end, an extension of the Neural Additive Model for competing risks settings is proposed in this paper, resulting in an inherently interpretable survival model. Our approach retains the interpretability and feature-wise transparency of NAMs and allowing for flexible, non-linear modelling of cause-specific or sub-distribution hazards in competing risks scenarios.
@@ 3.2 Risk-Specific Projections
For each feature $i$ and competing risk $k$, a separate linear projection transforms the feature representation to its contribution to the log-hazard ratio. To address scale ambiguities across different competing risks and ensure fair comparison of feature contributions, we constrain the projection vectors to have unit L2 norm.
The risk-specific projection is defined as:
$g_{i,k}(\mathbf{h}_{i})=\tilde{\mathbf{w}}_{i,k}^{T}\mathbf{h}_{i}$ (7)
where $\tilde{\mathbf{w}}_{i,k}$ is the L2-normalized projection vector:
$\tilde{\mathbf{w}}_{i,k}=\frac{\mathbf{w}_{i,k}}{\|\mathbf{w}_{i,k}\|_{2}+\epsilon}$ (8)
with $\mathbf{w}_{i,k}\in\mathbb{R}^{d}$ being the learnable weight vector for projection $i\in\{1,2,\ldots,p\}$ and risk $k\in\{1,2,\ldots,K\}$, and $\mathbf{h}_{i}\in\mathbb{R}^{d}$ being the feature representation.
This normalization constraint ensures that $\|\tilde{\mathbf{w}}_{i,k}\|_{2}=1$ for all feature-risk pairs, which constraints all projection vectors to operate on the same scale and enabling direct comparison of feature importance across different competing risks.
@@ 3.5 Partial Likelihood Loss Function
To train the model, the standard Cox partial likelihood approach, adapted for competing risks Prentice et al. (1978), is implemented. For a dataset with $N$ subjects, the negative log partial likelihood for event type $k\in\{1,\dots,K\}$ is
$\small\mathcal{L}_{k}=-\sum_{\begin{subarray}{c}n=1\\ E_{n}=k\end{subarray}}^{N}\Bigg[\eta_{k}(\mathbf{x}_{n})-\log\left(\sum_{\begin{subarray}{c}j=1\\ T_{j}\geq T_{n}\end{subarray}}^{N}\exp(\eta_{k}(\mathbf{x}_{j}))\right)\Bigg]$ (11)
where $E_{n}$ is the event type for subject $n$ and $T_{n}$ is their event or censoring time. The risk set at time $T_{n}$ consists of all subjects $j$ who have not yet experienced any event ($T_{j}\geq T_{n}$), and $\mathbf{x}_{j}$ denotes the feature vector for each subject $j$ in this risk set. The overall loss is the sum of the negative log partial likelihoods across all event types:
@@
where $\gamma$ is the $L_{2}$ regularization parameter and $\Theta$ represents all model parameters. Since many real-world problems involving competing risks suffer from class imbalance, we adopt a risk-frequency-weighted version of the partial likelihood. Specifically, we define:
$\small\mathcal{L}_{k,\omega}=-\omega_{k}\sum_{\begin{subarray}{c}n=1\\ E_{n}=k\end{subarray}}^{N}\bigg[\eta_{k}(\mathbf{x}_{n})-\log\Big(\sum_{\begin{subarray}{c}j=1\\ T_{j}\geq T_{n}\end{subarray}}^{N}\exp\big(\eta_{k}(\mathbf{x}_{j})\big)\Big)\bigg]$ (13)
@@
$\mathcal{L}_{\text{weighted}}=\sum_{k=1}^{K}\mathcal{L}_{k,\omega}$ (14)
@@ 3.6 Baseline Hazard Estimation
In the cause-specific proportional hazards formulation, the hazard for event type $k\in\{1,\dots,K\}$ at time $t\geq 0$ for a subject with covariates $\mathbf{x}\in\mathbb{R}^{p}$ is expressed as
$\lambda_{k}(t\mid\mathbf{x})=\lambda_{0k}(t)\cdot\exp\!\big(\eta_{k}(\mathbf{x})\big),$ (15)
where
  • $\lambda_{0k}(t)$ is the baseline cause-specific hazard function for event type $k$, independent of covariates,
  • $\eta_{k}(\mathbf{x})$ is the covariate-dependent log-risk function produced by the model.
We do not directly parameterize $\lambda_{0k}(t)$. Instead, we estimate the corresponding baseline cumulative hazard function $\lambda_{0k}(t)=\int_{0}^{t}\lambda_{0k}(u)\,du$ after training using the Breslow estimator Breslow (1975):
$\hat{\lambda}_{0k}(t_{m})=\sum_{\begin{subarray}{c}n:T_{n}\leq t\\ E_{n}=k\end{subarray}}\frac{1}{\sum_{j:T_{j}\geq T_{i}}\exp\!\left(\eta_{k}(\mathbf{x}_{j})\right)},$ (16)
where $T_{n}$ denotes the observed time for subject $n$, $E_{n}\in\{0,1,\dots,K\}$ is the event indicator ($E_{n}=0$ if censored), and the denominator represents the sum of relative risks for all subjects still at risk at time $T_{n}$.
This estimated baseline cumulative hazard $\hat{\lambda}_{0k}(t_{m})$ can then be used to compute the cumulative incidence function (CIF) for each event type $k$ at test time.
Estimating baseline hazards is necessary for several reasons. First, while the neural additive component of CRISP-NAM efficiently learns relative risks between subjects (hazard ratios), baseline hazard estimation enables translation of these relative measures into absolute risk predictions. This is required for clinical decision-making, where probabilities of events are needed. Second, in competing risks settings, accurate baseline hazard estimation is required for proper calculation of CIFs, as shown in Eqs. (17) and (18). Third, the baseline hazard captures the underlying temporal pattern of risk independent of covariates, allowing CRISP-NAM to generate time-dependent predictions at clinically relevant horizons (e.g., 1-year, 5-year risks). Finally, proper evaluation metrics such as Brier scores and time-dependent AUCs at specific time points depend on accurate absolute risk estimation.
@@ 3.7 Prediction of Absolute Risks
To predict the cumulative incidence function (CIF) (Gray, 1988) for each competing event, we use the relationship between cause-specific hazards and the CIF. For a subject with covariates $\mathbf{x}$, the CIF for event type $k$ at time $t\geq 0$ is defined as
$F_{k}(t|\mathbf{x})=\int_{0}^{t}S(u|\mathbf{x})\,\lambda_{k}(u|\mathbf{x})\,du,$ (17)
@@
  • $S(t|\mathbf{x})$ is the overall survival function, i.e., the probability of not experiencing any event up to time $t$,
  • $u$ is the integration variable representing time between 0 and $t$,
  • $\lambda_{k}(u|\mathbf{x})=\lambda_{0k}(u)\exp(\eta_{k}(\mathbf{x}))$ is the cause-specific hazard for cause $k$.
The survival function is given by
$S(t|\mathbf{x})=\exp\!\left(-\sum_{k=1}^{K}\int_{0}^{t}\lambda_{k}(u|\mathbf{x})\,du\right).$ (18)
In practice, a discrete approximation is used to compute these integrals. Let $\{t_{1},t_{2},\dots,t_{M}\}$ denote a set of ordered discrete time points with $t_{m}\in[0,t]$. Then the CIF can be approximated as
$\hat{F}_{k}(t|\mathbf{x})\approx\sum_{t_{m}\leq t}\hat{S}(t_{m-1}|\mathbf{x})\cdot\hat{\lambda}_{k}(t_{m}|\mathbf{x}),$ (19)
where
$\hat{S}(t_{m-1}|\mathbf{x})$ $=\exp\!\left(-\sum_{k^{\prime}=1}^{K}\sum_{t_{\ell}<t_{m}}\hat{\lambda}_{k^{\prime}}(t_{\ell}|\mathbf{x})\right),$
$\hat{\lambda}_{k}(t_{m}|\mathbf{x})$ $=\hat{\lambda}_{0k}(t_{m})\exp\!\big(\eta_{k}(\mathbf{x})\big),$
and $\hat{\lambda}_{0k}(t_{m})$ denotes the estimated baseline cause-specific hazard at time $t_{m}$ for event type $k$.